OpenStudy (anonymous):

ules kicks a soccer ball off the ground and into the air with an initial velocity of 25 feet per second. Assume the starting height of the ball is 0 feet. Approximately, how long does it take until the soccer ball hits the ground again? 0.6 sec 0.8 sec 1.6 secs 2.8 secs

OpenStudy (anonymous):

U=Uo*t-1/2g*t^2 where U=velocity, Uo=initial velocity, g=acceleration of gravity, 10m/s^2, and t=time. For U=0 you have 0=25t+1/2*g*t^2. Solve for t, and you get the ascension time . Multiply by 2 and you have the whole time.

OpenStudy (anonymous):

warning! G is 10m/s^2. To use it change feet into meters or meters into feet.